OpenStudy (anonymous):

3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

A

OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 7x - one over two (8x + 4) = 6 Step 1: 7x - 4x - 2 = 6 Step 2: 3x - 2 = 6 Step 3: 3x - 2 + 2 = 6 Step 4: 3x = 6 Step 5: x = 2 Which step was the error made and why? Step 1, because incorrect signs have been used Step 2, because the terms have been added incorrectly Step 3, because a number has been added only to one side of the equation No error was made

OpenStudy (anonymous):

step 3

OpenStudy (anonymous):

because if you were to add 2 to both sides, it would be : 3x - 2 + 2 = 6 + 2

OpenStudy (anonymous):

thank you :3 Jacob has four cabin bags. He wants to determine which cabin bag is the heaviest. The actual mass of the cabin bags is shown below. Bag A B C D Mass (in kg) 5.65 5.59 5.63 5.60 If the weighing machine measures to the nearest tenth of a kg, which bag will not be the same mass as the others? Bag A Bag B Bag C Bag D

OpenStudy (anonymous):

Bag A